Dustin's Door Company make and sell custom-made doors for families. Dustin's Door Company has provided its contribution format income statement for July.

Sales (6,200 units) $ 217,000
Variable expenses 117,800
Contribution margin 99,200
Fixed expenses 87,900
Net operating income $ 11,300

If Dustin's Door Company sells 6,400 units, its net operating income should be closest to: (Do not round intermediate calculations.)

Garrison 16e Rechecks 2018-06-19

Multiple Choice

  • $18,300

  • $14,500

  • $11,665

  • $11,300
Answer #1

Contribution margin per unit = 99,200/6200 = 16 per unit

Net operating income for 6400 units

= Total Contribution margin - Fixed expenses

= (6400*16) - 87,900

= 14,500